### v2

# Covering in Hamming and Grassmann Spaces: New Bounds and Reed–Solomon-Based Constructions

Abstract
We study covering problems in Hamming and Grassmann spaces through a unified coding-theoretic and information-theoretic framework. Viewing covering as a form of quantization in general metric spaces, we introduce the notion of the average covering radius as a natural measure of average distortion, complementing the classical worst-case covering radius. By leveraging tools from one-shot rate-distortion theory, we derive explicit non-asymptotic random-coding bounds on the average covering radius in both spaces, which serve as fundamental performance benchmarks.
On the construction side, we develop efficient puncturing-based covering algorithms for generalized Reed–Solomon (GRS) codes in the Hamming space and extend them to a new family of subspace codes, termed character-Reed–Solomon (CRS) codes, for Grassmannian quantization under the chordal distance. Our results reveal that, despite poor worst-case covering guarantees, these structured codes exhibit strong average covering performance. In particular, numerical results in the Hamming space demonstrate that RS-based constructions often outperform random codebooks in terms of average covering radius. In the one-dimensional Grassmann space, we numerically show that CRS codes over prime fields asymptotically achieve average covering radii within a constant factor of the random-coding bound in the high-rate regime. Together, these results provide new insights into the role of algebraic structure in covering problems and high-dimensional quantization.
Index Terms:
Covering problem, covering algorithm, Reed–Solomon codes, subspace codes.
I Introduction
Covering problems lie at the heart of several questions in coding theory, information theory, and signal processing, where one seeks structured families of objects that approximate the ambient space with minimal distortion. They are mathematically appealing in their own right, but have also found a wide range of technical applications in data compression, signal processing, clustering, sampling, and robust system design, especially in high-dimensional spaces, to name a few [37, 38]. In discrete settings, such as the Hamming space, the covering radius of a code captures its ability to represent arbitrary vectors, while in geometric settings like the Grassmann space, analogous notions arise through metrics such as the chordal distance on subspaces. Despite their shared conceptual foundation, these two domains have typically been treated with separate analytic tools and constructions. This work develops a unified perspective on covering in Hamming and Grassmann spaces, focusing on both the algorithmic task of finding nearby codewords and the quantitative task of understanding the distortion inherent to different code families.
The notion of subspace codes and their application in non-coherent communication was first developed in the seminal work of Koetter and Kschischang [18] in the context of randomized network coding. Specifically, they developed a coding theory framework for subspace coding over finite fields. Inspired by this, a parallel framework in the field of real/complex numbers was later developed by Soleymani and Mahdavifar [36]. They introduced a new algebraic construction for one-dimensional complex subspace codes called character-polynomial (CP) codes and later extended it to higher dimensions [35].
In the Hamming space, we study the covering problem through the lens of Reed–Solomon (RS) codes [27]. This well-known family of codes is not only theoretically rich but also widely used in practice, ranging from hard-disk drives to distributed storage and computing. Although various properties of RS codes have been studied extensively in the literature, they continue to inspire the development of new codes to this day. This includes the aforementioned CP codes [36], which are essentially the concatenation of a subcode of generalized Reed–Solomon (GRS) code, where a certain subset of the coefficients of the message polynomial is set to zeros, with a character function mapping the codeword coordinates to the complex unit circle. In a recent work, Gooty et al. [14] used CP codes to construct new precoding designs for multiple-input multiple-output (MIMO) systems with limited feedback, an important problem in wireless communications.
I-A Problem Motivation
Covering can be viewed as a form of quantization in a general metric space, which is central to problems involving the approximation or representation of continuous (or large discrete) spaces using structured, finite subsets. In a canonical covering problem, one seeks to cover a metric space with the minimum number of spheres of a prescribed radius, referred to as the covering radius. This guarantees that any point in the space can be approximated by mapping it to the center of a nearby sphere within the covering radius.
A natural approach to the covering problem is to leverage well-established code families in the underlying metric space that possess strong minimum-distance properties, while shifting attention to the algorithmic challenge of quantization. Specifically, the goal is to design efficient procedures that map an arbitrary element of the space to a nearby codeword, without requiring optimal (nearest-neighbor) decoding. In the Hamming space, GRS codes are known to meet the redundancy bound [17, Corollary 11.1.3], implying that their worst-case covering radius is large and therefore unattractive from a classical covering perspective. To the best of our knowledge, this remains the only general covering result available for RS codes.
Despite this unfavorable worst-case guarantee, we show in this paper that the average covering radius of GRS codes is significantly smaller than the redundancy bound. This observation highlights a fundamental gap between worst-case and typical covering behavior and provides strong motivation for developing efficient covering algorithms for RS-based constructions. Moreover, since subspace codes can be obtained by mapping RS codes to the complex domain, these insights naturally extend to covering problems in Grassmann spaces, with direct relevance to applications such as limited-feedback wireless communication systems.
I-B Related Work
The theory of covering codes has a rich history in classical information theory. Early work in this area includes characterizing upper and lower bounds on the covering radius and its computational complexity [5, 24], bounds on the minimal size of a code of a given covering radius [6, 39], the complexity of bounding the covering radius of a binary code [7], the relationship between the covering radius of a code and its subcodes [4], the covering radius of extended Reed–Solomon codes [10, 9], among others. We refer the reader to the extensive list of references following [17, Chapter 15]. More recently, there has been a renewed interest in determining the covering radius of various classes of codes, including Melas codes [31], Zetterberg type codes [32, 33], Euclidean codes [34], and extended algebraic-geometry codes [40], as well as bounds on the expected Hamming distortion of linear codes over $\mathbb{F}_{2}$ and lattices in $\mathbb{R}^{n}$ [25]. However, none of these recent works studied the problem of designing efficient covering algorithms and only focused on determining the covering radius. Related works in the subspace domain include asymptotic bounds and constructions [2, 3], covering bounds in Grassmann spaces over finite fields [26], efficient decoding algorithms for CP codes [28] and, more recently, sensing subspace codes [21, 13].
Despite the extensive early literature and the recent renewed interest in determining the covering radius of specific code families, the design and analysis of efficient covering algorithms has remained largely unexplored. Moreover, existing works predominantly focus on worst-case covering guarantees, leaving the typical or average-case behavior of covering schemes less understood. We hope that the present work helps renew interest in the covering problem by shifting attention toward algorithmic and average-case perspectives, particularly in light of its growing relevance to modern applications such as quantization, lossy compression, limited-feedback communication systems, and high-dimensional signal representation.
I-C Our Contribution
In this paper, we develop a unified algorithmic and information-theoretic framework for covering in both Hamming and Grassmann spaces. On the algorithmic side, we propose an efficient puncturing-based covering algorithm for GRS codes. The algorithm operates by successively puncturing the code and invoking an off-the-shelf decoder on the resulting shortened codes, until a nearby codeword is identified. We show that this procedure efficiently, i.e., with polynomial-time complexity, finds, for any input vector, a codeword within the covering radius of the input vector. We analyze the average-case performance of the proposed algorithm by characterizing the average number of punctures required for success and by estimating the resulting average covering radius, defined as the expected distance between the input vector and the selected codeword. In the Hamming space, we derive theoretical bounds on the average number of punctures and partially resolve a conjecture posed in the preliminary version of this work [29, Conjecture 1]. These results demonstrate that, despite their poor worst-case covering guarantees, GRS codes exhibit favorable average covering behavior.
To extend these ideas to subspace covering in the complex domain, we introduce character-Reed–Solomon (CRS) codes, which generalize CP codes by removing structural constraints tailored to packing. We show how the proposed puncturing-based algorithm can be adapted to CRS codes, yielding an efficient covering method for one-dimensional Grassmannian spaces over complex numbers under the chordal distance.
From an information-theoretic perspective, we formalize the notion of average covering radius as a natural measure of average distortion when a code is used for quantization. Building on one-shot rate-distortion results of Elkayam and Feder [11], we derive explicit non-asymptotic random-coding bounds on the average covering radius in both Hamming and Grassmann spaces. A numerical analysis of CRS codes over prime fields suggests that, at extreme high rates, these algebraic constructions asymptotically achieve average covering radii within a constant factor of the random-coding bound in the one-dimensional Grassmann space. Moreover, numerical results in the Hamming space demonstrate that RS-based codes frequently outperform random codebooks with respect to the average covering radius.
The rest of the paper is organized as follows. In Section II we provide the necessary background on GRS and subspace codes. In Section III we define the average covering radius and explore its connection to the average distortion measure in lossy source coding. In Section IV we present the covering algorithm for GRS codes along with a theoretical analysis of its performance as well as numerical results. In Section V we introduce CRS subspace codes and present a covering algorithm for these codes. Finally, we conclude the paper in Section VI with remarks on possible future directions.
II Preliminaries
Throughout this paper, $q$ denotes a positive integer power of a prime number $p$. Given a code ${\mathcal{C}}\subseteq{\mathbb{F}}_{q}^{n}$ and a codeword $c\in{\mathcal{C}}$, we denote the Hamming sphere of radius $\tau$ centered at $c$ by $B(c,\tau)$ and its volume by $\operatorname{Vol}_{q}(\tau,n):=|B(c,\tau)|$.
II-A Reed–Solomon Code
Fix $k\leq n\leq q$. Consider the message space:
${\mathcal{F}}(k,q)$ $:=\{f\in{\mathbb{F}}_{q}[X]:\deg(f)<k\},$
which consists of all polynomials of degree less than $k$ over ${\mathbb{F}}_{q}$. The elements of ${\mathcal{F}}(k,q)$ are called message polynomials, whose coefficients represent message symbols. The coefficient of $X^{j}$ in $f\in{\mathcal{F}}(k,q)$ is denoted by $f_{j}$.
Given distinct $\alpha_{1},\dots,\alpha_{n}\in{\mathbb{F}}_{q}$ (called evaluation points), the encoding of $f\in{\mathcal{F}}(k,q)$ in the Reed–Solomon code $\operatorname{RS}:=\operatorname{RS}_{n}({\mathcal{F}}(k,q))$ of length $n$ and dimension $k$ over ${\mathbb{F}}_{q}$ is defined as follows:
$\operatorname{RS}(f)$ $:=(f(\alpha_{1}),\dots,f(\alpha_{n})).$ (1)
In addition, given arbitrary $v_{1},\dots,v_{n}\in{\mathbb{F}}_{q}^{\times}$ (called column multipliers), the encoding of $f\in{\mathcal{F}}(k,q)$ in the generalized Reed–Solomon code $\operatorname{GRS}:=\operatorname{GRS}_{n}({\mathcal{F}}(k,q))$ of length $n$ and dimension $k$ over ${\mathbb{F}}_{q}$ is defined as follows:
$\operatorname{GRS}(f)$ $:=(v_{1}f(\alpha_{1}),\dots,v_{n}f(\alpha_{n})).$ (2)
$\operatorname{RS}$ and $\operatorname{GRS}$ above are well known to be maximum distance separable (MDS), i.e., they are $n,k,d]_{q}$ codes with $d:=n-k+1\geq 1$. Note also that $\operatorname{GRS}=\operatorname{RS}$ when $v_{i}=1$ for all $i$.
II-B Chordal Distance and Subspace Code
For an ambient vector space $W$, we use ${\mathcal{P}}(W)$ and ${\mathcal{P}}_{m}(W)$ to denote the set of all subspaces of $W$ and the set of all $m$-dimensional subspaces of $W$, respectively. In particular, ${\mathcal{P}}_{m}({\mathbb{C}}^{n})$ is referred to as a Grassmann space and is denoted by $G_{m,n}({\mathbb{C}})$. Any $U\in G_{m,n}({\mathbb{C}})$ is equipped with the natural inner product $\langle u,v\rangle:={u}^{\dagger}v$ for $u,v\in U$.
Definition 1 ([8, 36, § 3]).
Given $U,V\in G_{m,n}({\mathbb{C}})$, let $u_{i}\in U$ and $v_{i}\in V$ for $i\in\{1,\dots,m\}$ such that $|\langle u_{i},v_{i}\rangle|$ is maximal, subject to the condition that they form orthonormal bases for $U$ and $V$, respectively. The $i$-th principal angle $\theta_{i}$ between $U$ and $V$ is defined as $\theta_{i}:=\arccos|\langle u_{i},v_{i}\rangle|$. Then, the chordal distance between $U$ and $V$ is defined as follows:
$d_{{\rm ch}}(U,V):=\sqrt{\sum_{i=1}^{m}\sin^{2}\theta_{i}}.$ (3)
We define a subspace code (also called Grassmann code) to be a collection ${\mathcal{C}}\subseteq G_{m,n}({\mathbb{C}})$ of subspaces with respect to the chordal distance.
Soleymani and Mahdavifar [36] introduced a new class of subspace codes called character-polynomial (CP) codes defined as follows. First, we recall the definition of a character.
Definition 2 (Character).
A homomorphism $\chi$ from the additive group of ${\mathbb{F}}_{q}$ to the complex unit circle $U$ is called an (additive) character of ${\mathbb{F}}_{q}$. For $\beta\in{\mathbb{F}}_{q}$, the character $\chi_{\beta}$ of ${\mathbb{F}}_{q}$ is defined as follows:
$\chi_{\beta}:{\mathbb{F}}_{q}\to U,\quad\alpha\mapsto\exp\left(\frac{2\pi i\operatorname{Tr}(\beta\alpha)}{p}\right),$ (4)
where $\operatorname{Tr}:{\mathbb{F}}_{q}\to{\mathbb{F}}_{p}$ is the absolute trace function. It is well-known that every character $\chi$ of ${\mathbb{F}}_{q}$ is of the form $\chi=\chi_{\beta}$ for some $\beta\in{\mathbb{F}}_{q}$. The character $\chi_{0}$ is called the trivial character.
Definition 3 (CP Code [36, Definition 6]).
Fix $k\leq n<q$, a non-trivial character $\chi_{\beta}$ of ${\mathbb{F}}_{q}$, and units $\alpha_{1},\dots,\alpha_{n}\in{\mathbb{F}}_{q}^{\times}:={\mathbb{F}}_{q}\setminus\{0\}$. Define ${\mathcal{F}}_{p}(k,q)$ to be the set of all polynomials $f\in{\mathcal{F}}(k,q)$ with $f_{jp}=0$ for all integers $j\geq 0$. Then, the encoding of $f\in{\mathcal{F}}_{p}(k,q)$ in $\operatorname{CP}:=\operatorname{CP}_{n,\beta}({\mathcal{F}}_{p}(k,q))\subseteq G_{1,n}({\mathbb{C}})$ is defined as follows:
$\operatorname{CP}(f)$ $:=(\chi_{\beta}(f(\alpha_{1})),\dots,\chi_{\beta}(f(\alpha_{n}))),$ (5)
where $\operatorname{CP}(f)$ is identified with the one-dimensional subspace $\langle\operatorname{CP}(f)\rangle$.
Note that the structure of the message space ${\mathcal{F}}_{p}(k,q)$ is crucial for orthogonality and distance properties that make CP codes useful for packing subspaces in $G_{1,n}({\mathbb{C}})$ [36, 35]. Recall that in the sphere packing problem, the goal is to fit in the space pairwise disjoint spheres centered at the codewords. This is fundamentally related to decoding, where, given a point in the space, the objective is to uniquely determine its nearest codeword. In particular, given a packing of the space where the spheres have radius $\tau<d/2$, where $d$ is the minimum distance of the code, one can uniquely decode any message from $\tau$ errors, and the goal is to maximize the decoding radius $\tau$. The dual problem to sphere packing is sphere covering, which is the subject of our discussion in this work. In the sphere covering problem, the goal is to cover the space using minimally overlapping spheres. This is fundamentally related to quantization, which is the process of approximating any given point in the space by its closest codeword.
III Covering Radius as a Measure of Distortion
III-A Average Covering Radius
The (Hamming) covering radius of a block code ${\mathcal{C}}\subseteq{\mathbb{F}}_{q}^{n}$ is defined as follows [17]:
$\rho_{{\rm H}}({\mathcal{C}})$ $:=\max_{y\in{\mathbb{F}}_{q}^{n}}\min_{c\in{\mathcal{C}}}d_{{\rm H}}(y,c),$ (6)
where $d_{{\rm H}}$ denotes Hamming distance. For instance, it is well known that $\rho_{{\rm H}}(\operatorname{GRS})=d-1$. This means, in particular, that any $y\in{\mathbb{F}}_{q}^{n}$ is at distance at most d1d-1 from a $\operatorname{GRS}$ codeword.
Given a vector $y\in{\mathbb{F}}_{q}^{n}$ and a code ${\mathcal{C}}\subseteq{\mathbb{F}}_{q}^{n}$, the covering problem asks to find a codeword $c\in{\mathcal{C}}$ such that $d_{{\rm H}}(y,c)\leq\rho_{{\rm H}}({\mathcal{C}})$. An algorithm that solves the covering problem for $\mathcal{C}$} is called a covering algorithm for $\mathcal{C}$}.
Generally speaking, the covering radius of a code represents the maximum error when the code is used for the quantization of its ambient space. This is because the covering radius characterizes the worst-case scenario of the quantization process. In practice, however, given a certain distribution over the space, the average quantization error becomes more relevant. This motivates us to modify (6) to the following definition.
Definition 4 (Average Covering Radius).
For a block code ${\mathcal{C}}\subseteq{\mathbb{F}}_{q}^{n}$ and a distribution $\mathcal{D}$} on ${\mathbb{F}}_{q}^{n}$, the average (Hamming) covering radius of $\mathcal{C}$} over $\mathcal{D}$} is defined as follows:
$\overline{\rho}_{{\rm H}}({\mathcal{C}})$ $:=\operatorname*{\mathbb{E}}_{X\sim{\mathcal{D}}}\left[\min_{c\in{\mathcal{C}}}d_{{\rm H}}(X,c)\right].$ (7)
Similarly, for a subspace code ${\mathcal{C}}\subseteq G_{m,n}({\mathbb{C}})$ and a distribution $\mathcal{D}$} on $G_{m,n}({\mathbb{C}})$, the average (chordal) covering radius of $\mathcal{C}$} over $\mathcal{D}$} is defined as follows:
$\overline{\rho}_{{\rm ch}}({\mathcal{C}})$ $:=\operatorname*{\mathbb{E}}_{U\sim{\mathcal{D}}}\left[\min_{V\in{\mathcal{C}}}d_{{\rm ch}}(U,V)\right].$ (8)
Note that Definition 4 can be generalized naturally to codes over other metric spaces, e.g., rank-metric codes. Henceforth, covering radius will be understood to mean Hamming covering radius for the Hamming space, and chordal covering radius for the Grassmann space.
III-B Average Distortion of Random Codes
In the single-shot approach to the channel coding (or lossy compression) problem, the goal is to control a random object by providing achievable bounds without any assumption on the object itself. In particular, let $\mathcal{X}$} be a set of input symbols and $\mathcal{Y}$} a set of reproduction symbols. Given a code ${\mathcal{C}}\subseteq{\mathcal{Y}}$ and distribution ${\mathcal{D}}_{{\mathcal{X}}}$ on $\mathcal{X}$}, Elkayam and Feder [11] defined the average distortion associated with $\mathcal{C}$} as follows:
$D({\mathcal{C}})$ $:=\operatorname*{\mathbb{E}}_{X\sim{\mathcal{D}}_{{\mathcal{X}}}}\left[\min_{c\in{\mathcal{C}}}d(X,c)\right],$ (9)
where $d:{\mathcal{X}}\times{\mathcal{Y}}\to{\mathbb{R}}^{+}$ is the distortion function. They gave the following explicit formula for the average distortion of a random code ${\mathcal{C}}_{{\mathcal{Y}}}:=\{Y_{j}\sim{\mathcal{D}}_{{\mathcal{Y}}}:1\leq j\leq M\}$ consisting of $M$ codewords $Y_{j}$ drawn independently from a distribution ${\mathcal{D}}_{{\mathcal{Y}}}$ on ${\mathcal{P}}({\mathcal{Y}})$ [11, Theorem 2]:
$\hskip-5.69054pt\operatorname*{\mathbb{E}}_{{\mathcal{C}}_{{\mathcal{Y}}}}[D({\mathcal{C}}_{{\mathcal{Y}}})]=\int_{0}^{1}\tilde{D}(w,{\mathcal{D}}_{{\mathcal{Y}}})G_{M}^{\prime}(w)\,dw,$ (10)
where
$\hskip-14.22636pt\tilde{D}(w,{\mathcal{D}}_{{\mathcal{Y}}}):=$ (11)
$\hskip 43.36464pt\frac{1}{w}\cdot\operatorname*{\mathbb{E}}_{(X,Y)\sim{\mathcal{D}}_{{\mathcal{X}}}\times{\mathcal{D}}_{{\mathcal{Y}}}}[d(X,Y)\cdot\mathbf{1}_{\{p_{X,Y,U}\leq w\}}],$
$\hskip-14.22636ptp_{x,y,U}:=\Pr_{Y\sim{\mathcal{D}}_{{\mathcal{Y}}}}[d(x,Y)<d(x,y)]$ (12)
$\hskip 117.07924pt+U\cdot\Pr_{Y\sim{\mathcal{D}}_{{\mathcal{Y}}}}[d(x,Y)=d(x,y)],$
$\hskip-14.22636ptG_{M}(w):=-(1-w)^{M-1}\cdot[1+(M-1)w],$ (13)
and $U$ is a uniform random variable in $[0,1]$. Henceforth, to simplify notation, we will usually omit the distribution under $\Pr$ or $\operatorname*{\mathbb{E}}$ when it is clear from the random variable argument(s).
The distance function of a code serves as a natural measure of distortion when the code is used for the quantization of its ambient space. Under this identification, the average distortion $D({\mathcal{C}})$, defined in (9), of a code $\mathcal{C}$} coincides with the average covering radius of $\mathcal{C}$} defined in Definition 4, i.e., $\overline{\rho}_{{\rm H}}({\mathcal{C}})$ in (7) for the Hamming space (${\mathcal{X}}={\mathcal{Y}}={\mathbb{F}}_{q}^{n}$, $d=d_{{\rm H}}$), and $\overline{\rho}_{{\rm ch}}({\mathcal{C}})$ in (8) for the Grassmann space (${\mathcal{X}}={\mathcal{Y}}=G_{1,n}({\mathbb{C}})$, $d=d_{{\rm ch}}$). In particular, we can use (10) to estimate the average covering radius of random codes in both spaces as shown by the next two theorems.
Theorem 1.
Let ${\mathcal{D}}_{{\mathcal{Y}}}$ denote the uniform distribution on the Hamming space ${\mathcal{Y}}={\mathbb{F}}_{q}^{n}$. Then, the expected average covering radius of a random code ${\mathcal{C}}_{{\mathcal{Y}}}$ consisting of $M$ codewords drawn independently from ${\mathcal{D}}_{{\mathcal{Y}}}$ is given by:
$\operatorname*{\mathbb{E}}_{{\mathcal{C}}_{{\mathcal{Y}}}}[\overline{\rho}_{{\rm H}}({\mathcal{C}}_{{\mathcal{Y}}})]=$ (14)
$\quad M(M-1)\sum_{j=1}^{n}\bigg[j\left(A_{M,j}-A_{M-1,j}\right)+A_{M-1,j}\sum_{t=0}^{j-1}V_{t}\bigg],$
where
$V_{j}:=\frac{\operatorname{Vol}_{q}(j,n)}{q^{n}},$ (15)
$A_{i,j}:=\frac{(1-V_{j})^{i}-(1-V_{j-1})^{i}}{i}.$ (16)
Proof.
Note that $\Pr[d_{{\rm H}}(x,Y)\leq j]=V_{j}$ for any $x\in{\mathbb{F}}_{q}^{n}$ by (15). Hence, by (12),
$p_{x,y,U}$ $=V_{t-1}+U(V_{t}-V_{t-1}),$ (17)
where $t=d_{{\rm H}}(x,y)$ and $U$ is uniform in $[0,1]$.
For $(X,Y)\sim{\mathcal{D}}_{{\mathcal{X}}}\times{\mathcal{D}}_{{\mathcal{Y}}}$, where ${\mathcal{D}}_{{\mathcal{X}}}$ is any distribution on ${\mathcal{X}}={\mathcal{Y}}$, let $E_{t}$ denote the event $d_{{\rm H}}(X,Y)=t$. Then, by (11),
$\tilde{D}(w,{\mathcal{D}}_{{\mathcal{Y}}})$ $=\frac{1}{w}\sum_{t=1}^{n}t\Pr[d_{{\rm H}}(X,Y)\cdot{\bf 1}_{\{p_{X,Y,U}\leq w\}}=t]$
$=\frac{1}{w}\sum_{t=1}^{n}t\Pr[E_{t}\hbox{ and }p_{X,Y,U}\leq w]$
$=\frac{1}{w}\sum_{t=1}^{n}t\Pr[E_{t}]\cdot\Pr[p_{X,Y,U}\leq w|E_{t}].$ (18)
Now, by (15),
$\hskip-5.69054pt\Pr[E_{t}]$ $=\operatorname*{\mathbb{E}}[\Pr[E_{t}|X]]=\operatorname*{\mathbb{E}}[V_{t}-V_{t-1}]=V_{t}-V_{t-1}$ (19)
and, by (17),
$\hskip-5.69054pt\Pr[p_{X,Y,U}\leq w|E_{t}]$ $=\Pr\left[U\leq\frac{w-V_{t-1}}{V_{t}-V_{t-1}}\bigg|E_{t}\right]$
$=\begin{cases}0&\hbox{if }w\leq V_{t-1}\\ \dfrac{w-V_{t-1}}{V_{t}-V_{t-1}}&\hbox{if }V_{t-1}<w\leq V_{t}\\ 1&\hbox{if }w>V_{t}.\end{cases}$ (20)
Substituting (19) and (20) into (18) yields:
$\tilde{D}(w,{\mathcal{D}}_{{\mathcal{Y}}})=t_{0}\left(1-\frac{V_{t_{0}-1}}{w}\right)+\sum_{t=1}^{t_{0}-1}\frac{t}{w}(V_{t}-V_{t-1}),$ (21)
where $t_{0}:=t_{0}(w)\in\{1,\dots,n\}$ such that $V_{t_{0}-1}<w\leq V_{t_{0}}$. Substituting (21) and $G^{\prime}(w)=M(M-1)w(1-w)^{M-2}dw$ into (10) gives:
$\operatorname*{\mathbb{E}}_{{\mathcal{C}}_{{\mathcal{Y}}}}[\overline{\rho}_{{\rm H}}({\mathcal{C}}_{{\mathcal{Y}}})]=\int_{0}^{1}\tilde{D}(w,{\mathcal{D}}_{{\mathcal{Y}}})G^{\prime}(w)dw$
$=M(M-1)\sum_{j=1}^{n}\bigg[j\int_{V_{j-1}}^{V_{j}}\bigg(1-\frac{V_{j-1}}{w}\bigg)w(1-w)^{M-2}dw$
$\hskip 68.28644pt+\sum_{t=1}^{j-1}t(V_{t}-V_{t-1})\int_{V_{j-1}}^{V_{j}}(1-w)^{M-2}dw\bigg]$
$=M(M-1)\sum_{j=1}^{n}\bigg[j\underbrace{\int_{V_{j-1}}^{V_{j}}w(1-w)^{M-2}dw}_{A_{M,j}-A_{M-1,j}}$
$\hskip 14.22636pt-\bigg(jV_{j-1}-\sum_{t=1}^{j-1}t(V_{t}-V_{t-1})\bigg)\underbrace{\int_{V_{j-1}}^{V_{j}}(1-w)^{M-2}dw}_{A_{M-1,j}}\bigg]$
$=M(M-1)\sum_{j=1}^{n}\bigg[j\left(A_{M,j}-A_{M-1,j}\right)+A_{M-1,j}\sum_{t=0}^{j-1}V_{t}\bigg],$
where we have used (13) and (16). ∎
To obtain the analog of Theorem 1 for the Grassmann space ${\mathcal{Y}}=G_{1,n}({\mathbb{C}})$, we first introduce the following notation. For $V=\langle v\rangle\in{\mathcal{Y}}$ with $\left\|{v}\right\|=1$, the orthogonal projection operator onto $V$ is denoted by $P_{V}:=vv^{\dagger}$. Some key observations are:
  • For any $x,y\in G_{1,n}({\mathbb{C}})$,
    $d_{{\rm ch}}(x,y)$ $=\sqrt{1-\operatorname{tr}(P_{x}P_{y})}.$ (22)
  • $p_{x,y,U}=\Pr[d_{{\rm ch}}(x,Y)<d_{{\rm ch}}(x,y)]$ for any continuous distribution ${\mathcal{D}}_{{\mathcal{Y}}}$ by (12), which is independent of $U$, and
  • $\operatorname{tr}(P_{x}P_{Y}){}=\operatorname{tr}(P_{Qx}P_{QY})=|\hat{Y}_{1}|^{2}\sim{\rm Beta}(1,n-1)$ by unitary invariance, for any fixed $x:=\langle\hat{x}\rangle\in{\mathcal{Y}}$, $Y$ Haar-uniform on $\mathcal{Y}$}, and unitary $Q$ with $Q\hat{x}=\hat{e}_{1}$, so that the distributions of $Y$ and $QY:=\langle\hat{Y}\rangle$ are equal [19].
We then have the following result.
Theorem 2.
Let ${\mathcal{D}}_{{\mathcal{Y}}}$ denote the uniform distribution on the Grassmann space ${\mathcal{Y}}=G_{1,n}({\mathbb{C}})$. Then, the expected average covering radius of a random code ${\mathcal{C}}_{{\mathcal{Y}}}$ consisting of $M$ codewords drawn independently from ${\mathcal{D}}_{{\mathcal{Y}}}$ is given by:
$\operatorname*{\mathbb{E}}_{{\mathcal{C}}_{{\mathcal{Y}}}}[\overline{\rho}_{{\rm ch}}({\mathcal{C}}_{{\mathcal{Y}}})]=\\ \frac{2(n-1)}{2n-1}M(M-1)\operatorname{B}\left(1+\frac{2n-1}{2(n-1)},M-1\right),$ (23)
where $\operatorname{B}(\cdot,\cdot)$ is the beta function.
Proof.
For $(X,Y)\sim{\mathcal{D}}_{{\mathcal{X}}}\times{\mathcal{D}}_{{\mathcal{Y}}}$, where ${\mathcal{D}}_{{\mathcal{X}}}$ is any distribution on ${\mathcal{X}}={\mathcal{Y}}$, let $Z:=\operatorname{tr}(P_{X}P_{Y}){}$. Then, for any $x\in{\mathcal{X}}$, we have $Z\mid X=x\sim{\rm Beta}(1,n-1)$, so that $Z\sim{\rm Beta}(1,n-1)$. Therefore, by (22),
$p_{x,y,U}$ $=\Pr[d_{{\rm ch}}(X,Y)<t]$
$=\Pr[Z>1-t^{2}]$
$=t^{2(n-1)},$ (24)
where $t=d_{{\rm ch}}(x,y)$. Letting $T:=d_{{\rm ch}}(X,Y){}$ therefore gives:
$\tilde{D}(w,{\mathcal{D}}_{{\mathcal{Y}}})$ $=\frac{1}{w}\operatorname*{\mathbb{E}}[d_{{\rm ch}}(X,Y)\cdot{\bf 1}_{\{p_{X,Y,U}\leq w\}}]$
$=\frac{1}{w}\operatorname*{\mathbb{E}}[T\cdot{\bf 1}_{\{T^{2(n-1)}\leq w\}}]$
$=\frac{1}{w}\int_{0}^{w^{\frac{1}{2(n-1)}}}tf_{T}(t)dt$
$=\frac{2(n-1)}{w}\int_{0}^{w^{\frac{1}{2(n-1)}}}t^{2n-2}dt$
$=\frac{2(n-1)}{2n-1}w^{\frac{1}{2(n-1)}},$ (25)
which follows using (24). Thus, by (10),
$\operatorname*{\mathbb{E}}_{{\mathcal{C}}_{{\mathcal{Y}}}}[\overline{\rho}_{{\rm ch}}({\mathcal{C}}_{{\mathcal{Y}}})]=\int_{0}^{1}\tilde{D}(w,{\mathcal{D}}_{{\mathcal{Y}}})G^{\prime}(w)dw$
$=\frac{2(n-1)}{2n-1}M(M-1)\int_{0}^{1}w^{1+\frac{1}{2(n-1)}}(1-w)^{M-2}dw$
$=\frac{2(n-1)}{2n-1}M(M-1)\operatorname{B}\left(1+\frac{2n-1}{2(n-1)},M-1\right),$
which follows using (13) and (25), as desired. ∎
It is important to note that the conclusions of Theorem 1 and Theorem 2 do not depend on the source distribution ${\mathcal{D}}_{{\mathcal{X}}}$ over which the average covering radius of the random code ${\mathcal{C}}_{{\mathcal{Y}}}$ is defined.
IV Covering Using GRS Codes
Decoding algorithms have been the subject of study for many decades, and several decoding algorithms for GRS codes are known [17, 16]. In contrast, to the best of our knowledge, no explicit covering algorithms are known for GRS codes. Note that covering involves solving a more intricate minimax problem, which is computationally harder due to the need to analyze the worst-case scenario over the entire space. Although decoding algorithms may be used for covering, their performance is bound to be subpar as spheres of radius $\tau<d/2$ do not cover the space well. Therefore, as the first algorithm of its kind to the best of our knowledge, we propose Algorithm 1 and analyze its performance in this section.
IV-A Algorithm
Given an $n,k,d]_{q}$ GRS code $\mathcal{C}$} and an input vector $y$ in its ambient space, the algorithm first tries to decode $y$. If the decoder is successful, then we have found the closest codeword to $y$. If not, it repeatedly (up to d1d-1 times) punctures the code at the last coordinate and tries to decode the corresponding substring of $y$ until the decoder is successful, when it finally returns the re-encoding of the decoder output in the original code. Note that in an actual implementation of the algorithm, it may be desirable to have the message polynomial returned directly instead of its re-encoding, as we will see in Section V-B.
Algorithm 1 $\operatorname{GRS-cover}({\mathcal{C}},y)$
[…]
2:A codeword $c\in{\mathcal{C}}$ with $d_{{\rm H}}(y,c)\leq d-1$ Available Functions: $\operatorname{GRS-decode}$: a $\operatorname{GRS}$ decoder
[…]
4:for $i=1,\dots,d-1$ do
5:  $f\leftarrow\operatorname{GRS-decode}({\mathcal{C}},y)$
6:  if $f$ is not null then
7:   return ${\mathcal{C}}_{0}(f)$
8:  else
9:   puncture $\mathcal{C}$} at coordinate ni+1n-i+1
10:   $y\leftarrow y[1..n-i]$
11:  end if
12:end for
Theorem 3 below proves the correctness of Algorithm 1.
Theorem 3.
Given an $n,k,d]_{q}$ GRS code $\mathcal{C}$} and $y\in{\mathbb{F}}_{q}^{n}$, $\operatorname{GRS-cover}({\mathcal{C}},y)$ returns a codeword $c\in{\mathcal{C}}$ with $d_{{\rm H}}(y,c)\leq d-1$.
Proof.
Puncturing an $[n,k]$ GRS code at any coordinate gives an [n1,k][n-1,k] GRS code. Denote by $y^{(i)}$ and ${\mathcal{C}}_{i}$ the values of $y$ and $\mathcal{C}$} on lines 10 and 9, respectively, at step $i$. Plainly, for any $i$,
$d_{{\rm H}}(y^{(i+1)},{\mathcal{C}}_{i+1}(f))$ $\leq d_{{\rm H}}(y^{(i)},{\mathcal{C}}_{i}(f)).$ (26)
If $\operatorname{GRS-decode}({\mathcal{C}}_{i},y^{(i)})$ is successful, then by repeated application of (26),
$d_{{\rm H}}(y^{(i)},{\mathcal{C}}_{i}(f))\leq d_{{\rm H}}(y,{\mathcal{C}}(f))\leq d-1.$
It remains to show that $\operatorname{GRS-decode}({\mathcal{C}}_{i},y^{(i)})$ does indeed succeed for some $i<d$. To see this, note that, for each $i$,
$d_{\min}({\mathcal{C}}_{i+1})=d_{\min}({\mathcal{C}}_{i})-1.$ (27)
Hence, $\operatorname{GRS-decode}({\mathcal{C}}_{i},y^{(i)})$ will succeed in at most d1d-1 steps, since $d_{\min}({\mathcal{C}}_{i})=1$ when i=d1i=d-1, at which point any $y^{(i)}$ becomes a valid codeword. ∎
We emphasize that Theorem 3 does not claim that Algorithm 1 returns the closest codeword to the input vector, but rather guarantees returning of a codeword within the covering radius of the code (there might be multiple codewords within the covering radius of the input vector). This distinction is essential: finding the closest codeword (i.e., solving the minimum-distance decoding problem) is NP-hard in general, whereas the proposed covering algorithm achieves its guarantee with polynomial-time complexity. A detailed analysis of the complexity of Algorithm 1 will be provided later in this section.
IV-B Improvement Using List Decoding
Consider the following modification of Algorithm 1, where $\operatorname{GRS-decode}$ is the Guruswami–Sudan list decoder (GS) [16] and returns the best candidate in the list (i.e., the codeword with the shortest Hamming distance from $y$). There are several advantages of doing this. First, it was shown by McEliece [22, 23] that GS almost always returns a list of size $1$, whereby implying that it is essentially a unique decoder. Next, the decoding radius of GS is given by:
$\tau_{\operatorname{GS}}$ $:=n-1-\lfloor\sqrt{(k-1)n}\rfloor.$ (28)
This implies, in particular, that Algorithm 1 will now return a $\operatorname{GRS}$ codeword $c$ within distance $\tau_{\operatorname{GS}}$ of $y$, i.e., with $d_{{\rm H}}(y,c)\leq\tau_{\operatorname{GS}}$, if such a codeword exists [23]. This potentially reduces the average number of punctures, thereby improving the average covering radius.
IV-C Complexity Analysis
Plainly, the worst-case running time of Algorithm 1 is $(n-k)\cdot T(n)$, where $T(n)$ is the running time of $\operatorname{GRS-decode}$.
  • For unique decoding, algorithms like Berlekamp–Welch (BW) are known with $T(n)=O(n^{3})$ [15, Theorem 15.1.4] as well as more efficient ones with $T(n)=O(n^{2})$ and $T(n)=O(n\operatorname{polylog}(n))$ [15, § 15.4].
  • For list decoding, there are known implementations of $\operatorname{GS}$ with $T(n)=O(s^{4}n^{2})$ [22, § VII], [22, § VII], [12], [30, § V], where $s>0$ is an integer parameter known as the interpolation multiplicity.
IV-D Average Number of Punctures
In practical settings, the average-case running time of an algorithm is generally of more interest rather than the worst-case. Note that the average-case running time of Algorithm 1 depends on the average number of punctures needed for $\operatorname{GRS-decode}$ to succeed. Based on the data in Table I, we made the following conjecture in an earlier version of this work [29].
Conjecture 1.
For an $n,k,d]_{q}$ GRS code, the average number of punctures needed for Algorithm 1 to succeed in returning a codeword within the covering radius is less than
  • $c_{1}(d-1)$ for $\operatorname{GRS-decode}$ a unique decoder, and
  • $c_{2}$ for $\operatorname{GRS-decode}$ the $\operatorname{GS}$ list decoder,
for some positive constants $c_{1}<1$ and $c_{2}$.
However, it turns out that the correct behavior of Algorithm 1 is given by the following theorem.
Theorem 4.
Suppose that $\operatorname{GRS-decode}$ in Algorithm 1 is a list decoder that returns all codewords within its decoding radius. Then, averaged uniformly on all inputs $y\in{\mathbb{F}}_{q}^{n}$, the number $N_{{\rm punc}}$ of punctures needed for $\operatorname{GRS-cover}({\mathcal{C}},y)$ to succeed for an $n,k,d]_{q}$ GRS code $\mathcal{C}$} satisfies the following inequalities:
$d-1-\frac{1}{q^{d-1}}\sum_{i=0}^{d-2}q^{i}\operatorname{Vol}_{q}(\tau_{i},n-i)\leq\overline{N}_{{\rm punc}}\\ \leq d-1-\frac{1}{q^{d-1}}\sum_{i=0}^{d-2}\frac{q^{i}\operatorname{Vol}_{q}(\tau_{i},n-i)}{L_{i}},$ (29)
where $\tau_{i}$ and $L_{i}$ are respectively the decoding radius and the maximum list size of $\operatorname{GRS-decode}({\mathcal{C}}_{i},\cdot)$.
Proof.
Using the notation from the proof of Theorem 3, let
$S_{i}$ $:=\{(y,c)\in{\mathbb{F}}_{q}^{n}\times{\mathcal{C}}_{i}:d_{{\rm H}}(y^{(i)},c)\leq\tau_{i}\},$ (30)
$\ell_{i}(y)$ $:=\#\{c\in{\mathcal{C}}_{i}:d_{{\rm H}}(y^{(i)},c)\leq\tau_{i}\}.$ (31)
In other words, $S_{i}$ is the set of all pairs $(y,c)$ such that the output list of $\operatorname{GRS-decode}({\mathcal{C}}_{i},y^{(i)})$ contains a polynomial $f$ with $c={\mathcal{C}}_{i}(f)$, and $\ell_{i}(y)$ is the size of the list. Then, the number of vectors $y\in{\mathbb{F}}_{q}^{n}$ for which Algorithm 1 succeeds within $i$ punctures is given by:
$V_{i}$ $:=\#\{y\in{\mathbb{F}}_{q}^{n}:\ell_{i}(y)\geq 1\}.$ (32)
In particular, the average number of punctures $\overline{N}_{{\rm punc}}$ needed for Algorithm 1 to succeed is given by:
$\frac{1}{q^{n}}\sum_{i=1}^{d-1}i(V_{i}-V_{i-1})$ $=\frac{1}{q^{n}}\left((d-1)V_{d-1}-\sum_{i=0}^{d-2}V_{i}\right)$
$=d-1-\frac{1}{q^{n}}\sum_{i=0}^{d-2}V_{i},$ (33)
since $V_{d-1}=q^{n}$. Now, by (30) and (31),
$|S_{i}|$ $=\sum_{y\in{\mathbb{F}}_{q}^{n}}\ell_{i}(y)=\sum_{c\in{\mathcal{C}}_{i}}q^{i}|B(c,\tau_{i})|$
$=q^{k+i}\operatorname{Vol}_{q}(\tau_{i},n-i).$ (34)
Since $L_{i}$ is the maximum list size of $\operatorname{GRS-decode}({\mathcal{C}}_{i},\cdot)$,
$1_{\{\ell_{i}(y)\geq 1\}}$ $\leq\ell_{i}(y)\leq 1_{\{\ell_{i}(y)\geq 1\}}L_{i}.$ (35)
Summing (35) over $y\in{\mathbb{F}}_{q}^{n}$ using (32) and (34) gives
$V_{i}$ $\leq|S_{i}|\leq V_{i}L_{i},$ (36)
i.e.,
$\frac{|S_{i}|}{L_{i}}$ $\leq V_{i}\leq|S_{i}|.$ (37)
Finally, (29) follows from (IV-D), (34) and (37). ∎
If $\operatorname{GRS-decode}$ is a unique decoder, we can substitute $L_{i}=1$ for all $i$ into Theorem 4 and obtain the following corollary.
Corollary 1.
Suppose that $\operatorname{GRS-decode}$ in Algorithm 1 is a unique decoder with maximum decoding radius. Then, averaged uniformly over all inputs $y\in{\mathbb{F}}_{q}^{n}$, the number $N_{{\rm punc}}$ of punctures needed for $\operatorname{GRS-cover}({\mathcal{C}},y)$ to succeed for an $n,k,d]_{q}$ GRS code $\mathcal{C}$} is given by:
$\overline{N}_{{\rm punc}}$ $=d-1-\frac{1}{q^{d-1}}\sum_{i=0}^{d-2}q^{i}\operatorname{Vol}_{q}(\tfrac{n-k-i}{2},n-i).$ (38)
In particular,
$(d-1)\left(1-\frac{1}{q}\right)$ $\leq\overline{N}_{{\rm punc}}\leq d-1,$ (39)
i.e., $\overline{N}_{{\rm punc}}/(d-1)\to 1$ as $q\to\infty$.
Proof.
Taking $\tau_{i}=(n-k-i)/2$ and $L_{i}=1$ for all $i$ in (29) gives (38), and in (36) gives, by (34),
$V_{i}$ $=|S_{i}|=q^{k+i}\operatorname{Vol}_{q}(\tau_{i},n-i).$ (40)
Thus,
$\frac{1}{q^{n}}\sum_{i=0}^{d-2}V_{i}$ $\leq\frac{(d-1)V_{d-2}}{q^{n}}$
$\leq\frac{(d-1)q^{k+d-2}\operatorname{Vol}_{q}(\tau_{d-2},n-d+2)}{q^{n}}$
$=\frac{(d-1)\operatorname{Vol}_{q}(\tfrac{1}{2},k+1)}{q}$
$=\frac{d-1}{q},$ (41)
since the $V_{i}$ are non-decreasing. Substituting (41) into (IV-D) gives (39). ∎
Corollary 1 clearly shows that the first part of Conjecture 1 does not hold in general. We believe that a similar analysis that carefully takes into account intersections of Hamming spheres (e.g., see Theorem 5 below) can also help resolve the second part of the conjecture.
IV-E Coverage Fraction
Next, we provide a lower bound on the fraction of space covered by Hamming spheres of a given radius centered at the codewords of an MDS code in terms of its weight and intersection distribution. The key observation is that one can explicitly compute the sizes of the Hamming spheres and their pairwise intersections, as Theorem 5 below demonstrates.
Theorem 5.
Let $\mathcal{C}$} be an $n,k,d]_{q}$ MDS code. Then, the fraction of the space ${\mathbb{F}}_{q}^{n}$ covered by Hamming spheres of radius $\tau$ associated with $\mathcal{C}$} is lower bounded as follows:
$\frac{1}{q^{n}}{\raisebox{-0.25pt}{$\biggl\lvert$}}{\bigcup_{c\in{\mathcal{C}}}B(c,\tau)}\raisebox{-0.25pt}{$\biggr\rvert$}\\ \geq\frac{1}{q^{d-1}}\left(\sum_{i=0}^{\tau}\binom{n}{i}(q-1)^{i}-\frac{1}{2}\sum_{w=d}^{2\tau}A_{w}I(w,\tau)\right),$ (42)
where the weight distribution $A_{w}$ is given by:
$A_{w}:=\binom{n}{w}\sum_{j=0}^{w-d}(-1)^{j}\binom{w}{j}(q^{w-d+1-j}-1),$ (43)
and the intersection distribution $I(w,\tau)$ is given by:
$I(w,\tau):=$ (44)
$\sum_{z=0}^{n-w}\binom{n-w}{z}(q-1)^{n-w-z}\hskip-8.53581pt\sum_{\begin{subarray}{c}n-\tau-z\leq u,v\leq\tau\\ u+v\leq w\end{subarray}}\binom{w}{u}\binom{w-u}{v}(q-2)^{w-u-v}$.
Proof.
By the inclusion-exclusion principle,
${\raisebox{-0.25pt}{$\biggl\lvert$}}{\bigcup_{c\in{\mathcal{C}}}B(c,\tau)}\raisebox{-0.25pt}{$\biggr\rvert$}\geq\sum_{c\in{\mathcal{C}}}|B(c,\tau)|-\frac{1}{2}\sum_{c_{1}\neq c_{2}}|B(c_{1},\tau)\cap B(c_{2},\tau)|\\ =|{\mathcal{C}}|\left(\sum_{i=0}^{\tau}\binom{n}{i}(q-1)^{i}-\frac{1}{2}\sum_{w=d}^{2\tau}A_{w}I(w,\tau)\right),$ (45)
where $A_{w}$ is the number of codewords $c\in{\mathcal{C}}$ of weight $w$, and $I(w,\tau):=|B(c_{1},\tau)\cap B(c_{2},\tau)|$ for $c_{1},c_{2}\in{\mathcal{C}}$ with $d_{{\rm H}}(c_{1},c_{2})=w$. Now, (43) is known to hold for $d\leq w\leq n$ (e.g., see [20]), and (44) follows by a counting argument, where $u$ (resp. $v$) represents the number of indices in $\{1,\dots,n\}$ where $y$ and $c_{1}$ (resp. $c_{2}$) agree, and $z$ represents the number of indices where $c_{1}$, $c_{2}$ and $y$ agree. Substituting $|{\mathcal{C}}|=q^{k}=q^{n-d+1}$ into (45) gives (42). ∎
Note that Theorem 5 is sharp when overlaps of three or more Hamming spheres are negligible, which is the case when $\tau$ is sufficiently small, and, in particular, expected to be the case for the smallest $\tau$ such that the union of Hamming spheres of radius $\tau$ covers most of the space. We observe numerically in Section IV-F3 that the radius $\tau_{\max}$ yielding the best lower bound is generally close to $\tau_{\operatorname{GS}}$ defined in (28).
IV-F Simulation Results and Comparison with Random Coding Bound
IV-F1 Fixed Length
Table I lists the empirical average number of punctures for $\operatorname{GRS-cover}$ to succeed on an $n,k]_{q}$ GRS code $\mathcal{C}$} with $\operatorname{GRS-decode}$ a unique decoder (BW) and a list decoder (GS), for $(q,n)\in\{(7,6),(11,10),(17,16)\}$, and $1\leq k<n$. For each $k$, the average was computed over $500$ Monte Carlo simulations. It shows that the number of punctures required for Algorithm 1 to succeed using BW is close to the worst-case covering radius of the code, but is much smaller when using the GS list decoder. In other words, while Hamming spheres of radius <d/2<d/2 do not cover the space well, those of radius $\tau_{\operatorname{GS}}$ cover a large fraction of the space.
$k$ BW GS
1 3.836 0
2 2.44 0.006
3 1.654 0.08
4 0.532 0.264
5 0.868 0
$q=7$, $n=6$
$k$ BW GS
1 8.224 0
2 6.872 0.452
3 5.654 0.366
4 4.34 0.378
5 3.022 0.7
6 1.864 0.744
7 1.428 0.036
8 0.376 0.16
9 0.922 0
$q=11$, $n=10$
$k$ BW GS
1 14.573 0
2 13.293 0.69
3 12.253 1.65
4 10.927 0.95
5 9.847 2.097
6 8.493 0.837
7 7.22 0.61
8 6.18 1.027
9 4.843 0.76
10 3.66 1.173
11 2.717 0.2
12 1.587 0.627
13 1.187 0.007
14 0.18 0.073
15 0.947 0
$q=17$, $n=16$
TABLE I: Average number of punctures before $\operatorname{GRS-cover}$ succeeds using unique (BW) and list (GS) decoders
We also estimated the average covering radius $\overline{\rho}_{{\rm H}}({\mathcal{C}})$ of $\mathcal{C}$} by computing the average value of $d_{{\rm H}}(y,c)$ over $500$ simulations for $(q,n)=(7,6)$ and $1\leq k<n$ with $\operatorname{GRS-decode}$ a unique decoder (BW), a list decoder (GS), and a MAP decoder, where $c$ is the output codeword of $\operatorname{GRS-cover}({\mathcal{C}},y)$. The results are compared in Figure 1 with the worst-case covering radius $\rho_{{\rm H}}({\mathcal{C}})=n-k$ and the random coding bound (14) given by Theorem 1. It can be seen that Algorithm 1 using GS list decoder improves over random codes for all $k$ (despite the fact that GRS codes are not ideal for covering) and the resulting average covering radius is close to that obtained using the MAP decoder. In fact, Algorithm 1 using BW also improves over random codes for sufficiently large $k$. The decoding radius parameter $\tau$ of GS in each case was taken to be its maximum decoding radius $\tau_{\operatorname{GS}}$ defined in (28). Note that the MAP decoder provides the optimal covering solution with an exponential complexity.
11223344551122334455Dimension ($k$)Average Covering RadiusUpper Bound (nkn-k) Random Codes (Theorem 1)Algorithm 1 (BW)Algorithm 1 (GS)Algorithm 1 (MAP)
Figure 1: Comparison of simulated average covering radii of a $6,k]_{7}$ GRS code and random codes for $1\leq k<6$
IV-F2 Fixed Rate
4(a) shows the average covering radius of a $n,k]_{q}$ GRS code $\mathcal{C}$}, where n=q1n=q-1 and $k=\lfloor nR\rfloor$, for $5\leq q<32$ given by Algorithm 1 with $\operatorname{GRS-decode}=\operatorname{GS}$ as well as the average number of punctures required by the algorithm to succeed, both computed over $300$ Monte Carlo simulations for $R\in\{1/3,1/2,2/3\}$. The list size parameter of $\operatorname{GS}$ in each case was taken to be the length of the code. This suggests that the average number of punctures and the average covering radius both increase with the covering radius $\rho_{{\rm H}}({\mathcal{C}})=n-k$ when the rate $R$ of the code is fixed and the decoding radius of $\operatorname{GS}$ is sub-optimal.
(a) Simulated average covering radius of a $q-1,\lfloor(q-1)R\rfloor]_{q}$ GRS code given by $\operatorname{GRS-cover}$ with $\operatorname{GRS-decode}=\operatorname{GS}$ and number of punctures until it succeeds for $5\leq q<32$
IV-F3 $\tau_{\operatorname{GS}}$ vs. $\tau_{\max}$
Figure 4 shows the fraction of space covered by Hamming spheres of radius close to $\tau_{\operatorname{GS}}$, and that $\tau_{\operatorname{GS}}=\tau_{\max}$ for $(q,n,k)=(17,14,2)$. Figure 5 shows a comparison of $\tau_{\operatorname{GS}}$ with $\tau_{\max}$ for $(q,n)=(47,46)$ and $1\leq k<n$. Both support our hypothesis that Hamming spheres of radius $\tau_{\operatorname{GS}}$ cover almost all of the space.
Figure 4: Lower bound on fraction of space covered by Hamming spheres of radius $\tau\in(d/2,d)$ for a $14,2,d]_{17}$ GRS code
Figure 5: $\tau_{\max}$ vs. $\tau_{\operatorname{GS}}$ for $q=47,n=46$ and $1\leq k\leq 37$
V Character-Reed–Solomon Code
Recall from Definition 3 that the message space ${\mathcal{F}}_{p}(k,q)$ of CP codes consists of polynomials $f\in{\mathcal{F}}(k,q)$ with $f_{jp}=0$ for all integers $j\geq 0$. This condition was used to establish minimum distance properties of CP codes [36]. Since our focus in this work is on covering rather than packing, we utilize the full message space ${\mathcal{F}}(k,q)$ and generalize the definition of CP codes below to what we call character-Reed–Solomon (CRS) codes.
Definition 5 (CRS Code).
Fix $k\leq n<q$, a non-trivial character $\chi_{\beta}$ of ${\mathbb{F}}_{q}$, and distinct evaluation points $\alpha_{1},\dots,\alpha_{n}\in{\mathbb{F}}_{q}$. Then, the encoding of $f\in{\mathcal{F}}(k,q)$ in the CRS code $\operatorname{CRS}:=\operatorname{CRS}_{n,\beta}({\mathcal{F}}(k,q))\subseteq G_{1,n}({\mathbb{C}})$ is given by:
$\operatorname{CRS}(f)$ $:=(\chi_{\beta}(f(\alpha_{1})),\dots,\chi_{\beta}(f(\alpha_{n}))),$ (46)
where we identify $\operatorname{CRS}(f)$ with the one-dimensional subspace $\langle\operatorname{CRS}(f)\rangle$.
We now derive some theoretical properties of CRS codes, including bounds on the code size.
V-A Theoretical Properties and Bounds on Code Size
Note that $\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,q))$ is essentially the concatenation of a RS code with the character map $\chi_{\beta}:{\mathbb{F}}_{q}\to U$, and therefore can be regarded as a RS-coded modulation. In particular, $\operatorname{CRS}_{n,\beta}({\mathcal{F}}(k,q))$ prior to concatenation by $\chi_{\beta}$ is simply expressed as $\operatorname{RS}_{n}({\mathcal{F}}(k,q))$ using (1). Hence,
$|\operatorname{CRS}|$ $\leq|\operatorname{RS}|=|{\mathcal{F}}(k,q)|=q^{k}.$ (47)
Equality holds if and only if the encoding map (46) is injective. To see when this is the case, observe that $\operatorname{CRS}(f)=\operatorname{CRS}(g)$ for $f,g\in{\mathcal{F}}(k,q)$ if and only if $\chi_{\beta}(f(\alpha_{j}))=\chi_{\beta}(g(\alpha_{j}))$ for $j\in\{1,\dots,n\}$, which holds if and only if $\operatorname{Tr}(\beta h(\alpha_{j}))=0$ for each $j$ by (4), where $h=f-g\in{\mathcal{F}}(k,q)$. The following lemma characterizes the zeros of $\operatorname{Tr}$.
Lemma 1.
Let $\alpha\in{\mathbb{F}}_{q}$. Then, $\operatorname{Tr}(\alpha)=0$ if and only if
$\alpha$ $=\beta^{p}-\beta,$ (48)
for some $\beta\in{\mathbb{F}}_{q}$. If $\beta=\beta_{0}$ is a solution to (48), then all solutions are given by $\beta_{\lambda}=\beta_{0}+\lambda$ for $\lambda\in{\mathbb{F}}_{p}$.
Proof.
Consider the map:
$\phi:{\mathbb{F}}_{q}\to{\mathbb{F}}_{q},\quad\beta\mapsto\beta^{p}-\beta.$ (49)
Note that $\phi$ is ${\mathbb{F}}_{p}$-linear with $\ker(\phi)={\mathbb{F}}_{p}$ by Fermat’s little theorem. Hence, $|\phi({\mathbb{F}}_{q})|=q/p$ by the first isomorphism theorem. But ${\mathbb{F}}_{q}\supsetneq\ker(\operatorname{Tr})\supseteq\phi({\mathbb{F}}_{q})$, whence $\ker(\operatorname{Tr})=\phi({\mathbb{F}}_{q})$. Thus, if $\alpha\in\ker(\operatorname{Tr})$, then $\alpha=\phi(\beta_{\lambda})$ for some $\lambda\in{\mathbb{F}}_{p}$. ∎
Therefore, $\beta h(\alpha_{j})=\beta_{j}^{p}-\beta_{j}$ for some $\beta_{j}\in{\mathbb{F}}_{q}$, for each $j$. Interpolating $u\in{{\mathcal{F}}(n,q)}$ over $\{(\alpha_{j},\beta_{j}):1\leq j\leq n\}$ implies that $u^{p}-u-\beta h$ vanishes on the set $\{\alpha_{1},\dots,\alpha_{n}\}$. Furthermore, by Lemma 1, there are exactly $p^{n}$ such $u$, each corresponding to a realization of $(\beta_{1},\dots,\beta_{n})$
We thus obtain the following result.
Theorem 6.
Let ${\mathcal{C}}=\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,q))$ with evaluation points $\alpha_{1},\dots,\alpha_{n}$. Then, ${\mathcal{C}}(f)={\mathcal{C}}(g)$ if and only if
$\beta$ $(f-g)\equiv u^{p}-u{}\pmod{(X-\alpha_{1})\cdots(X-\alpha_{n})}$, (50)
for some $u\in{\mathcal{F}}(n,q)$. In particular, given the ${\mathbb{F}}_{p}$-linear map
$T:{\mathcal{F}}(k,q)\to{\mathbb{F}}_{p}^{n},\quad f\mapsto(\operatorname{Tr}(\beta f(\alpha_{j})){})_{j=1}^{n},$ (51)
we have
$|{\mathcal{C}}|$ $=|\operatorname{Im}(T)|=p^{\operatorname{rank}(T)}.$ (52)
Furthermore, if there exists a $u\in{\mathcal{F}}(n,q)$ satisfying (50), then there are exactly $p^{n}$ such $u$. Finally, if (50) is an equality, i.e.,
$\beta(f-g)$ $=u^{p}-u$ (53)
for some $u\in{\mathcal{F}}(k/p,q)$, then (53) holds for exactly $p$ such $u$. Otherwise, $\deg(u)\geq n/p$ for all $u$ satisfying (50).
Proof.
Only the claims regarding (53) remain to be proved.
Since $\deg(f-g)<k<n$, it follows from (50) that $\deg(u)\geq n/p$ or $\deg(u)<k/p$, with (50) taking the form of an equality precisely in the latter case.
Consider the map $\phi$ in (49) extended to ${\mathbb{F}}_{q}[X]$. Then $\phi$ is ${\mathbb{F}}_{p}$-linear with $\ker(\phi)={\mathbb{F}}_{p}$. Therefore, if $h\in\phi({\mathbb{F}}_{q}[X])$, i.e., $h=u^{p}-u$ for some $u\in{\mathbb{F}}_{q}[X]$, then there are exactly $p$ such $u\in{\mathbb{F}}_{q}[X]$, given by $u+\lambda$ for $\lambda\in{\mathbb{F}}_{p}$. ∎
Observe that (50) cannot be satisfied for $f\neq g$ when $k$ is sufficiently small. This is made precise in the corollary below.
Corollary 2.
Let ${\mathcal{C}}=\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,q))$. If $k<1+np/q$, then $|{\mathcal{C}}|=q^{k}$.
Proof.
We claim that, under the given hypothesis, the map $T$ given by (51) is injective. To see this, note that if $T(f)=0$, then the polynomial $\operatorname{Tr}\circ\beta f\in{\mathbb{F}}_{q}[X]$ has zeros at the $n$ evaluation points of $\mathcal{C}$}. But
$\deg(\operatorname{Tr}\circ\beta f)$ $=\deg(\operatorname{Tr})\deg(\beta f)\leq(k-1)q/p<n,$
which forces $\operatorname{Tr}\circ\beta f=0$, i.e. $f=0$, thereby proving the claim. Thus, $|{\mathcal{C}}|=p^{\dim_{{\mathbb{F}}_{p}}{\mathcal{F}}(k,q)}=q^{k}$ by (52), as desired. ∎
In general, Theorem 6 implies that $|{\mathcal{C}}|$ must be equal to a power of $p$ less than or equal to $p^{n}$. Combined with (47), this gives the following bounds.
Theorem 7.
Let ${\mathcal{C}}=\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,q))$. Then $|{\mathcal{C}}|$ is a power of $p$ such that
$\frac{p^{n}}{q^{n-k}}$ $\leq|{\mathcal{C}}|\leq\min\{p^{n},q^{k}\}.$ (54)
Equality holds on both sides of (54) when $k=n$ or $q=p$.
Proof.
Let $\alpha_{1},\dots,\alpha_{n}$ be the evaluation points of $\mathcal{C}$}, and put $w(X)=(X-\alpha_{1})\cdots(X-\alpha_{n})$. Let $R:={\mathbb{F}}_{q}[X]/(w)$ and consider the map $\phi$ in (49) extended to $R$. By the Chinese remainder theorem,
$R$ $\cong\prod_{i=1}^{n}{\mathbb{F}}_{q}[X]/(X-\alpha_{i})\cong{\mathbb{F}}_{q}^{n},$ (55)
so that $\ker(\phi)\cong{\mathbb{F}}_{p}^{n}$ by Lemma 1. Moreover, by Theorem 6,
${\mathcal{C}}(f)={\mathcal{C}}(g)$ $\iff\beta(f-g)\in S\cap\phi(R),$
where $S={\mathcal{F}}(k,q)/(w)\subseteq R$. Therefore, by the isomorphism theorems,
$|{\mathcal{C}}|$ $=\frac{|S|}{|S\cap\phi(R)|}\geq\frac{|S|}{|\phi(R)|}=\frac{q^{k}}{(q/p)^{n}},$ (56)
and
$|{\mathcal{C}}|$ $=\frac{|S+\phi(R)|}{|\phi(R)|}\leq\frac{|R|}{|\phi(R)|}=|\ker(\phi)|=p^{n}.$ (57)
Thus, (54) follows from (56), (57) and (47). Equality holds in (56) if and only if $\phi(R)\subseteq S$, which is satisfied for $k=n$ or $q=p$ (the latter since $u(\alpha_{i})^{q}=u(\alpha_{i})$ for all $i$). Likewise, equality holds in (57) if and only if $S+\phi(R)=R$, which is satisfied for sufficiently large $k$, in particular, for $k=n$. ∎
Note that if $k<n$, it is possible that neither $\phi(R)\subseteq S$ nor $S+\phi(R)=R$ holds. On the other hand, if $q=p$, then $\phi(R)=\{0\}$. Hence, $S+\phi(R)=R$ holds if and only if $k=n$.
V-B Covering Algorithm
As mentioned in the introduction, CP codes have recently been used for the quantization problem in MISO systems with limited feedback [14], and the quantization error was characterized using the covering radius of CP codes. Moreover, a decoding algorithm for CP codes over prime fields was recently proposed [28]. However, as discussed at the beginning of Section IV, decoding algorithms do not produce good coverage due to the decoding radius being much smaller than the covering radius. Therefore, we propose Algorithm 2 below for covering lines in $G_{1,n}({\mathbb{C}})$ using CRS codes. To describe the idea behind the algorithm, consider the map $\psi_{\beta}$ defined in (58), which can be regarded as an extension of the 11-to-$q/p$ map $\chi_{\beta}^{-1}$ to all of $\mathbb{C}$}, obtained by first mapping any $z\in{\mathbb{C}}$ to its closest point (in euclidean distance) in $\chi_{\beta}({\mathbb{F}}_{q})$ then taking its pre-image:
$\psi_{\beta}:{\mathbb{C}}$ $\to\binom{{\mathbb{F}}_{q}}{q/p}$ (58)
$z$ $\mapsto\chi_{\beta}^{-1}\left(\exp\left(\frac{2\pi i}{p}\left\lfloor\frac{p\arg(z)}{2\pi}+\frac{1}{2}\right\rfloor\right)\right).$
Here, $\binom{S}{t}$ conventionally denotes the set of all subsets of $S$ of size $t$. The given subspace $\langle y\rangle\in G_{1,n}({\mathbb{C}})$ is first transformed into a $\operatorname{RS}$ codeword by applying $\psi_{\beta}$ to each $y_{i}$ and picking a random element in $\chi_{\beta}^{-1}(y_{i})$, which is then followed by a call to $\operatorname{GRS-cover}$ on the resulting vector. Finally, the message polynomial corresponding to the output is encoded in the CRS code and returned by the algorithm.
Algorithm 2 $\operatorname{CRS-cover}({\mathcal{C}},y)$
1:CRS code ${\mathcal{C}}=\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,q))$, arbitrary subspace $\langle y\rangle\in G_{1,n}({\mathbb{C}})$
2:A codeword $c\in{\mathcal{C}}$ with $d_{{\rm ch}}(\langle y\rangle,\langle c\rangle)\leq\rho_{{\rm ch}}({\mathcal{C}})$ Available Functions: $\psi_{\beta}$, $\operatorname{GRS-cover}$, $\operatorname{GRS-decode}$
[…]
4:for $i=1,\dots,n$ do
5:   $u_{i}\leftarrow\hbox{random element in }\psi_{\beta}(y_{i})$
6:end for
[…]
9:return ${\mathcal{C}}(f)$
In a practical implementation of Algorithm 2, one may avoid Step 8 altogether by modifying $\operatorname{GRS-cover}$ in Algorithm 1 to directly return the message polynomial rather than its re-encoding in Step 7. Moreover, the pre-images $u_{j}\in\psi_{\beta}(y_{j})$ in Step 5 under the character map $\chi_{\beta}$ may be chosen using a “best-of-NN” approach for some fixed $N$. In other words, out of $N$ random assignments of the elements of $\mathbb{F}_{q}$ to the $u_{i}$ under $\psi_{\beta}$, the one giving the smallest chordal distance of the $\operatorname{CRS}$ codeword from $y$ is chosen. This is because, in general, each point in $\chi_{\beta}(\mathbb{F}_{q})$ has $q/p$ distinct pre-images in $\mathbb{F}_{q}$, so a brute-force approach would need to test $(q/p)^{n}$ assignments to find the optimal one.
V-C Improved Upper Bound on Average Covering Radius
Let $\Psi=\chi_{\beta}({\mathbb{F}}_{q})$. Then $\Psi$ consists of the $p$-th roots of unity. Given the input vector $y\in{\mathbb{C}}^{n}$, let $\theta_{j}=\arg(y_{j})\in[0,2\pi)$ and, for $j\in\{1,\dots,n\}$, define
$\check{y}_{j}$ $:=\exp\left(\frac{2\pi i}{p}\left\lfloor\frac{p\theta_{j}}{2\pi}+\frac{1}{2}\right\rfloor\right).$ (59)
In other words, $\check{y}_{j}$ is the closest point to $y_{j}$ in $\Psi$ (in Euclidean distance). The following theorem improves the upper bound on the average chordal covering radius of $\operatorname{CP}_{n,\beta}({\mathcal{F}}_{p}(k,p))=\operatorname{CRS}_{n,{\beta}}({\mathcal{F}}(k,p))$ by Gooty et al. [14, Theorem 1].
Theorem 8.
Consider a distribution $\mathcal{D}$} on ${\mathbb{C}}^{n}$, where the amplitudes are i.i.d. with mean $\mu$ and variance $\sigma^{2}$, and are independent of the phases. Then, over $\mathcal{D}$}, the average covering radius of ${\mathcal{C}}:=\operatorname{CRS}_{n,\beta}({\mathcal{F}}(k,p))$ of rate $R:=k/n$ such that
$R$ $\geq\frac{1}{c+1}+\frac{(1-c)\sigma^{2}}{2(1+c)n\mu^{2}},$ (60)
where $c:=\sqrt{\cos(2\pi/p)}$, is bounded as follows:
$\overline{\rho}_{{\rm ch}}({\mathcal{C}})\leq\min\bigg\{\sqrt{1-\dfrac{(cR+R-1)^{2}\mu^{2}}{\mu^{2}+\sigma^{2}}},$ (61)
[…]
Proof.
The first expression inside the $\min$ in (61) is the square root of the bound on the mean-squared quantization error given by Gooty et al. [14, Theorem 1]. To obtain the improved bound, let $f\in{\mathcal{F}}(k,p)$ denote the output of $\operatorname{GRS-decode}$ in Step 8 of $\operatorname{CRS-cover}({\mathcal{C}},y)$. Then, by the triangle inequality,
$|\langle y,{\mathcal{C}}(f)\rangle|$ $\geq{\raisebox{-0.25pt}{$\biggl\lvert$}}{{\raisebox{-0.25pt}{$\biggl\lvert$}}{\sum_{i\in{\mathcal{A}}}y_{i}^{*}\chi_{\beta}(f(\alpha_{i}))}\raisebox{-0.25pt}{$\biggr\rvert$}-{\raisebox{-0.25pt}{$\biggl\lvert$}}{\sum_{i\in{\mathcal{B}}}y_{i}^{*}\chi_{\beta}(f(\alpha_{i}))}\raisebox{-0.25pt}{$\biggr\rvert$}}\raisebox{-0.25pt}{$\biggr\rvert$}$
$\geq{\raisebox{-0.25pt}{$\biggl\lvert$}}{\sum_{i\in{\mathcal{A}}}y_{i}^{*}\chi_{\beta}(f(\alpha_{i}))}\raisebox{-0.25pt}{$\biggr\rvert$}-\sum_{i\in{\mathcal{B}}}|y_{i}|,$ (62)
where
$\begin{aligned} {\mathcal{A}}&=\{1\leq i\leq n:\check{y}_{i}=\chi_{\beta}(f(\alpha_{i}))\},\\ {\mathcal{B}}&=\{1\leq i\leq n:\check{y}_{i}\neq\chi_{\beta}(f(\alpha_{i}))\}.\end{aligned}$
To estimate the first term, note that $\check{y}_{j}=e^{i\delta_{j}}y_{j}/{|y_{j}|}$ for $j\in\{1,\dots,n\}$ by (59), where
$\delta_{j}$ $:=\frac{2\pi}{p}\left\lfloor\frac{p\theta_{j}}{2\pi}+\frac{1}{2}\right\rfloor-\theta_{j}\in\left[-\frac{\pi}{p},\frac{\pi}{p}\right),$ (63)
so that for $j\in{\mathcal{A}}$,
$y_{j}^{*}\chi_{\beta}(f(\alpha_{j}))$ $=y_{j}^{*}\check{y}_{j}=|y_{j}|e^{i\delta_{j}}.$ (64)
Thus, by (63) and (64),
$\hskip-20.00003pt{\raisebox{-0.25pt}{$\biggl\lvert$}}{\sum_{j\in{\mathcal{A}}}y_{j}^{*}\chi_{\beta}(f(\alpha_{j}))}\raisebox{-0.25pt}{$\biggr\rvert$}={\raisebox{-0.25pt}{$\biggl\lvert$}}{\sum_{j\in{\mathcal{A}}}|y_{j}|e^{i\delta_{j}}}\raisebox{-0.25pt}{$\biggr\rvert$}$
$=\sqrt{\sum_{j\in{\mathcal{A}}}|y_{j}|^{2}+2\sum_{\begin{subarray}{c}j,k\in{\mathcal{A}}\\ j<\ell\end{subarray}}|y_{j}||y_{\ell}|\cos(\delta_{j}-\delta_{\ell})}$
$\geq\sqrt{\sum_{j\in{\mathcal{A}}}|y_{j}|^{2}+2\sum_{\begin{subarray}{c}j,\ell\in{\mathcal{A}}\\ j<\ell\end{subarray}}|y_{j}||y_{\ell}|\cos(2\pi/p)}$
$\geq c\sum_{j\in{\mathcal{A}}}|y_{j}|.$
Let
${\mathcal{E}}$ $:=\bigg\{y\in{\mathcal{D}}:c\sum_{j\in{\mathcal{A}}}|y_{j}|\geq\sum_{j\in{\mathcal{B}}}|y_{j}|\bigg\}.$
By Markov’s inequality,
$\Pr[{\mathcal{E}}]$ $=\Pr\bigg[c\sum_{j=1}^{n}|y_{j}|\geq(c+1)\sum_{j\in{\mathcal{B}}}|y_{j}|\bigg]$
$\geq 1-\frac{\operatorname*{\mathbb{E}}\left[\sum_{j\in{\mathcal{B}}}|y_{j}|/\sum_{j=1}^{n}|y_{j}|\right]}{c/(c+1)}.$ (65)
Letting $A:=|{\mathcal{A}}|$, $B:=|{\mathcal{B}}|$, $Y:=\sum_{j=1}^{n}|y_{j}|$, and using the fact that the $|y_{j}|$ are i.i.d., we have
$\operatorname*{\mathbb{E}}\left[\frac{\sum_{j\in{\mathcal{B}}}|y_{j}|}{Y}\right]$ $=\operatorname*{\mathbb{E}}\left[\sum_{j=1}^{n}\frac{|y_{j}|}{Y}{\bf 1}_{\{j\in{\mathcal{B}}\}}\right]$
$=\sum_{j=1}^{n}\operatorname*{\mathbb{E}}\left[\frac{|y_{j}|}{Y}\right]\operatorname*{\mathbb{E}}[{\bf 1}_{\{j\in{\mathcal{B}}\}}]$
$=\frac{\operatorname*{\mathbb{E}}[B]}{n}$
$\leq\frac{n-k}{n}$
$=1-R,$ (66)
which follows by linearity and independence of the amplitudes and phases. Here we have used
$B=n-A\leq\rho_{{\rm H}}({\mathcal{C}})=\rho_{{\rm H}}(\operatorname{RS}_{n}({\mathcal{F}}(k,p))=n-k,$ (67)
which follows from the definition of $\rho_{{\rm H}}$ (see (6) and its associated remarks) and the injectivity of $\chi_{\beta}:{\mathbb{F}}_{p}\to{\mathbb{C}}$. Thus, by the law of total expectation,
$\operatorname*{\mathbb{E}}[|\langle y,{\mathcal{C}}(f)\rangle|^{2}]\geq\operatorname*{\mathbb{E}}[|\langle y,{\mathcal{C}}(f)\rangle|^{2}|{\mathcal{E}}]\Pr[{\mathcal{E}}]$
$\hskip-2.84526pt\geq\left(1-\frac{1-R}{c/(c+1)}\right)\cdot\operatorname*{\mathbb{E}}\left[\bigg(c\sum_{i\in{\mathcal{A}}}|y_{i}|-\sum_{i\in{\mathcal{B}}}|y_{i}|\bigg)^{2}\bigg|{\mathcal{E}}\right]$
$\hskip-2.84526pt=\frac{cR+R-1}{c}\cdot\operatorname*{\mathbb{E}}\left[\operatorname*{\mathbb{E}}\left[\bigg(c\sum_{i\in{\mathcal{A}}}|y_{i}|-\sum_{i\in{\mathcal{B}}}|y_{i}|\bigg)^{2}\bigg|{\mathcal{A}}\right]\bigg|{\mathcal{E}}\right]$
$\hskip-2.84526pt=\frac{cR+R-1}{c}\cdot\operatorname*{\mathbb{E}}[(cA-B)^{2}\mu^{2}+(c^{2}A+B)\sigma^{2}|{\mathcal{E}}]$
$\hskip-2.84526pt\geq\frac{cR+R-1}{c}\cdot[(ck+k-n)^{2}\mu^{2}+(c^{2}k+n-k)\sigma^{2}],$ (68)
which follows from (62), (65) and (66) upon noting that
  • $(cA-B)^{2}\mu^{2}+(c^{2}A+B)\sigma^{2}$ is increasing in $A$ for
    $A$ $\geq\frac{n}{1+c}+\frac{(1-c)\sigma^{2}}{2(1+c)\mu^{2}}.$ (69)
  • (69) is implied by (60), since $A\geq k=nR$ by (67).
Applying Jensen’s inequality twice followed by (68) gives:
$\overline{\rho}_{{\rm ch}}({\mathcal{C}})=\operatorname*{\mathbb{E}}\left[\sqrt{1-\frac{|\langle y,{\mathcal{C}}(f)\rangle|^{2}}{n\left\|{y}\right\|^{2}}}\right]$
$\leq\sqrt{1-\operatorname*{\mathbb{E}}\bigg[\frac{|\langle y,{\mathcal{C}}(f)\rangle|^{2}}{n\left\|{y}\right\|^{2}}\bigg]}$
$\leq\sqrt{1-\frac{cR+R-1}{c}\cdot\frac{(ck+k-n)^{2}\mu^{2}+(c^{2}k+n-k)\sigma^{2}}{n\operatorname*{\mathbb{E}}[\left\|{y}\right\|^{2}]}}.$
Substituting $\operatorname*{\mathbb{E}}[\left\|{y}\right\|^{2}]=n(\mu^{2}+\sigma^{2})$ yields the conclusion. ∎
Theorem 8 provides sufficient SNR and rate conditions under which CRS codes achieve a small average covering radius. Comparing this bound with the random-coding expectation of Theorem 2 yields the following corollary, which shows that CRS codes achieve average covering radii within a constant factor of random codes at high rate and high SNR.
Corollary 3.
Consider a distribution $\mathcal{D}$} on ${\mathbb{C}}^{n}$, where the amplitudes are i.i.d. with mean $\mu$ and variance $\sigma^{2}{}$, and independent of the phases, such that $\mu^{2}/\sigma^{2}\to\infty$ as $n\to\infty$. Then, for sufficiently large $n$, the average covering radius over $\mathcal{D}$} of a $\operatorname{CRS}$ code of rate $R>1-1/(\ln p)$ over ${\mathbb{F}}_{p}$ is within a factor of $2\sqrt{p(1-R)}$ of the average covering radius of a random code of size $p^{nR}$ in $G_{1,n}({\mathbb{C}})$.
Proof.
By Theorem 2, the expected average covering radius of a random code of size $M=p^{nR}$ is equal to
$\operatorname*{\mathbb{E}}_{{\mathcal{C}}_{{\mathcal{Y}}}}[\overline{\rho}_{{\rm ch}}({\mathcal{C}}_{{\mathcal{Y}}})]$ $=\frac{M(M-1)}{a_{n}-1}\operatorname{B}(a_{n},M-1)$
$=\frac{M(M-1)\Gamma(a_{n})\Gamma(M-1)}{(a_{n}-1)\Gamma(a_{n}+M-1)}$
$=\frac{\Gamma(a_{n}-1)\Gamma(M+1)}{\Gamma(a_{n}+M-1)},$ (70)
where and $\Gamma$ is the gamma function and
$a_{n}$ $:=2+\frac{1}{2(n-1)}.$ (71)
Using Stirling’s approximation, as $n\to\infty$,
$\frac{\Gamma(M+1)}{\Gamma(a_{n}+M-1)}$ $\sim\frac{M^{M+1/2}e^{-M}}{(a_{n}+M-2)^{a_{n}+M-3/2}e^{-(a_{n}+M-2)}}$
$=e^{a_{n}-2}\frac{M^{M+1/2}}{(a_{n}+M-2)^{a_{n}+M-3/2}}$
$=e^{a_{n}-2}\frac{M^{M+1/2-a_{n}-M+3/2}}{(\frac{a_{n}+M-2}{M})^{a_{n}+M-3/2}}$
$=\left(\frac{e}{M}\right)^{a_{n}-2}\frac{1}{(1+\frac{a_{n}-2}{M})^{a_{n}+M-3/2}}$
$\sim\left(\frac{e}{M}\right)^{a_{n}-2}\frac{1}{e^{a_{n}-2}(1+\frac{a_{n}-2}{M})^{a_{n}-5/2}}$
$\sim M^{2-a_{n}}$
$=M^{-\frac{1}{2(n-1)}}$
$\sim p^{-\frac{R}{2}},$ (72)
and, by Theorem 8,
$\overline{\rho}_{{\rm ch}}({\mathcal{C}})$ $\leq\sqrt{1-\frac{(cR+R-1)^{2}\mu^{2}}{\mu^{2}+\sigma^{2}}}$
$\sim\sqrt{1-(2R-1)^{2}},$ (73)
where ${\mathcal{C}}=\operatorname{CRS}_{n}({\mathcal{F}}_{p}(k,p))$. Here, the asymptotic notation $A(n)\sim B(n)$ conventionally means that $A(n)/B(n)\to 1$ as $n\to\infty$. Letting $R=1-\varepsilon$ for $\varepsilon<1/(\ln p)$, we have
$p^{-\frac{R}{2}}$ $=\sqrt{p^{-1}e^{\varepsilon\ln p}}\sim\sqrt{\frac{1+\varepsilon\ln p}{p}},$ (74)
and
$\sqrt{1-(2R-1)^{2}}$ $=\sqrt{1-(1-2\varepsilon)^{2}}\sim 2\sqrt{\varepsilon},$ (75)
i.e., the ratio of the average covering radii of a random code and $\mathcal{C}$}, by Eqs. 70, 71, 72, 73, 74 and 75, is asymptotically at least
$\frac{p^{-\frac{R}{2}}}{\sqrt{1-(2R-1)^{2}}}$ $\sim\frac{1}{2\sqrt{\varepsilon p}}.$ (76)
Therefore, $\mathcal{C}$} achieves the average covering radius of random codes within a factor of $2\sqrt{\varepsilon p}$ for sufficiently large $n$. ∎
Note that $2\sqrt{p(1-R)}\sim 2\sqrt{n-k}{}$ whenever $p=n+o(n)$. Classical results from analytic number theory (see, e.g., [1]) guarantee the existence of such $p$ for any $n$.
V-D Simulation Results and Comparison with Random Coding Bound
Figure 6 shows the Monte Carlo simulated average covering radius of ${\mathcal{C}}=\operatorname{CRS}_{6,1}({\mathcal{F}}(k,7))$ for $1\leq k<6$ computed using Algorithm 2 with $\operatorname{GRS-decode}\in\{\hbox{BW},\hbox{GS}\}$, together with theoretical bounds (compare this with Figure 1). The input vector $y\in{\mathbb{C}}^{n}$ for each simulation was chosen to have i.i.d. coordinates drawn from a complex normal distribution so that the amplitudes follow a Rayleigh distribution with mean $1$. The average chordal distance from $y$ to the output codeword was computed over $500$ data points for each $k$. The random coding bound is given by (23) in Theorem 2. The Gooty et al. upper bound is given by the square root of their bound on the mean squared quantization error [14, Theorem 1], which is improved upon by Theorem 8.
11223344550.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.90.911Dimension ($k$)Average Covering RadiusUpper Bound [14] Upper Bound (Theorem 8) Algorithm 2 (BW)Algorithm 2 (GS)Random Codes (Theorem 2)
Figure 6: Comparison of simulated average covering radii of a $(6,k)_{7}$ CRS code and random codes for $1\leq k<6$
The behavior observed in Figure 6 highlights an important distinction between covering in the Hamming and Grassmann spaces. While Reed–Solomon-based constructions can outperform random codes in the Hamming space under the average covering radius metric, the corresponding CRS constructions fall short of random codebooks in the Grassmann space. This performance gap can be attributed to a structural constraint inherent to CRS codes: all coordinates of the basis vector have equal amplitude, as defined in (46). Although this constant-modulus property is highly desirable in wireless communication applications, where it enables power-efficient transmission and simplifies hardware implementation, it restricts the geometric flexibility of the codebook in the Grassmann space. In contrast, random codebooks implicitly exploit variations in coordinate amplitudes, which allows for more effective space covering. We believe that achieving or surpassing the performance of non-asymptotic random codes in the Grassmann space requires the design of structured codes with non-uniform amplitude profiles across coordinates. The exploration of such amplitude-adaptive constructions, while preserving practical constraints, is left for future work.
Figure 7 compares the upper bound on the average covering radius of an $(n,k)_{p}$ CRS code and the random coding bound in the extreme high-rate regime $k/n\to 1$, where k=n1k=n-1 and $p$ is the smallest prime >n>n, for $n\leq 10^{4}$. The bounds are evaluated under a source distribution with mean $\mu=n$ and variance $\sigma^{2}=1$. As $n\to\infty$, both the random-coding bound and the upper bound derived in Theorem 8 vanish, while the ratio between them approaches $2$. In particular, the average covering radius of the CRS code remains sandwiched between the two bounds and within a fixed constant factor of the random-coding bound, in accordance with Corollary 3.
1,0001{,}0002,0002{,}0003,0003{,}0004,0004{,}0005,0005{,}0006,0006{,}0007,0007{,}0008,0008{,}0009,0009{,}00000.020.020.040.040.060.060.080.08Length ($n$)Average Covering RadiusUpper Bound (Theorem 8)Random Codes (Theorem 2)
Figure 7: Bounds on the average covering radius of an $(n,n-1)_{p}$ CRS code for $p=n+o(n)$
VI Conclusion and Future Directions
In this work, we developed efficient covering algorithms for both block and subspace codes based on RS codes. Central to our approach is a puncturing-based covering algorithm that leverages existing decoding techniques and enables efficient identification of codewords within the covering radius of the code. For block codes, we showed that when combined with the Guruswami–Sudan list decoder, the proposed algorithm achieves an average covering radius that closely approaches the optimal performance of the MAP decoder, despite operating with polynomial-time complexity.
From an information-theoretic perspective, we introduced the notion of the average covering radius for both block and subspace codes as a meaningful measure of average distortion in quantization problems. Using tools from one-shot compression and rate-distortion theory, we derived explicit random-coding bounds on the average covering radius in both Hamming and Grassmann spaces. These bounds provide fundamental benchmarks against which structured code constructions can be evaluated. In particular, our numerical results demonstrated that the RS-based constructions outperform random codebooks in various fixed block-length scenarios. The analysis provides new insights into the role of algebraic structures in covering problems and applications.
Several directions for future work naturally emerge from this study. First, it would be of interest to extend the present analysis to CP/CRS codes over arbitrary finite fields, as well as to more general Grassmannian code constructions. Second, the performance of the CRS covering algorithm may be further improved by refining the mapping step in Step 5 of Algorithm 2, potentially through more deterministic or structure-aware choices. Additionally, the upper bound in Theorem 8 could be tightened by replacing Jensen’s inequality with sharper concentration bounds or by incorporating additional sources of randomness into the analysis.
A notable limitation of CRS codes is the constant-modulus constraint on codeword coordinates. This, while advantageous for practical wireless implementations, restricts geometric flexibility and likely contributes to their inferior performance relative to random codes in short-length regimes. This observation motivates the exploration of multi-level CRS codes with variable coordinate amplitudes, and a systematic study of their impact on average covering performance.
Finally, many of the concepts and techniques developed in this work extend naturally beyond the Hamming and Grassmann spaces. In particular, investigating average covering properties and efficient covering algorithms for codes in other metric spaces, such as rank-metric codes, is another promising avenue for future research.